# Zhuosheng Zhang♣,†, Gongshen Liu♣,†

### v1

@@ Abstract
Pengzhou Cheng♣,thanks:   Equal contribution. Corresponding authors., Zheng Wu♣,∗, Zongru Wu, Aston Zhang, Zhuosheng Zhang♣,†, Gongshen Liu♣,† School of Electronic Information and Electrical Engineering, Shanghai Jiao Tong University {cpztsm520,wzh815918208,wuzongru,zhangzs,lgshen}@sjtu.edu.cn
@@ 1 Introduction
To annotate the confidence scores for the probed GUI agents in real-world scenarios, we first design a collaborative confidence probing framework. Inspired by Chen et al. (2024a), this framework integrates a layout parsing model Tang et al. (2019) and the most capable proprietary model, $\mathtt{GPT}$-$\mathtt{4o}$ Achiam et al. (2023) to function as a critic model. The critic model is used to supervise plan scheduling and confidence score based on our curated instructions to address complex scenarios. This framework is the first toolkit designed to identify when human intervention is necessary, generate confidence scores, and facilitate the automated construction of GUI trajectories. To further integrate confidence scoring into the probed GUI agent, we validate and refine these GUI trajectories and then fine-tune the model. This approach ensures action prediction accuracy while improving adaptability of human intervention.
@@ 2.1 MLLM-powered GUI Agents
Early efforts to build GUI agents rely on the availability of commercial MLLMs. These agents can be built through prompt learning based on $\mathtt{GPT}$-$\mathtt{4o}$ or Gemini-Pro Vision, e.g., AppAgent Zhang et al. (2023) and Mobile-Agent Wang et al. (2024a). However, practitioners are concerned about the costs associated with API requests and the delays in inference on mobile devices. Recent studies have focused on fine-tuning to optimize foundation models. On the one hand, they work on performing fine-grained visual understanding Bai et al. (2023), model scaling laws Chen et al. (2024b), multimodal information integration Hong et al. (2024), and GUI grounding enhancements Wu et al. (2024b); Qin et al. (2025) in the pre-training phase. On the other hand, researchers fine-tune the foundation model on GUI-specific datasets to enhance action orientation Wu et al. (2024a), planning decision Zhang et al. (2024c), perception enhancement Ma et al. (2024b), and reasoning Zhang and Zhang (2024); Zhang et al. (2024b). Moreover, a framework based on reinforcement learning (RL) designed specifically for the GUI agents can further enhance robustness Zhou et al. (2024); Liu et al. (2024); Wang et al. (2024b).
@@ 2.2 Capability Probing for GUI Agent
GUI agent-oriented capability probing is critical for real-world applications Deka et al. (2017). Generally, the capability of GUI agents can be probed by releasing benchmark datasets. Examples like UIBert Bai et al. (2021), SeeClick Cheng et al. (2024), and OS-Copilot Wu et al. (2024b), which investigate the problem of grounding understanding to UI elements on a screen. Besides, large-scale, diverse, and high-quality trajectory datasets can identify challenges of action prediction in terms of effectiveness (e.g., PixelHelp Li et al. (2020), Meta-GUI Sun et al. (2022), and AndroidWorld Rawles et al. (2024)), task complexity (e.g., Mobile-Bench Deng et al. (2024) and GUI Odyssey Lu et al. (2024)), and data-scaling (e.g., AITW Rawles et al. (2024) and AndroidControl Li et al. (2024)). After identifying the capability bottleneck of GUI agents, the introduction of specific strategies (e.g., planning lists Zhang et al. (2024c), action chains Zhang and Zhang (2024); Zhang et al. (2024b), and supplementary data) further enhance the environment perception. However, most benchmark datasets rely on crowdsourcing and human annotation.
@@ 3 Pilot Experiments
In this section, we first define GUI agent paradigms and then investigate the over-execution issue of the existing GUI agent on three complex scenarios. As shown in Figure 2, GUI agents confront substantial risks in real-world scenarios, such as those involving ambiguous user instructions (e.g., information absence and account logout), unexpected interruptions (e.g., hallucinations and shortcuts), and environmental hijacks (e.g., Pop-up hijacking and permission unauthorized).
@@ 3.1 GUI Agent Paradigm
Interactive GUI Agent. Given an interactive GUI agent $\mathcal{F}_{i}$ and system prompt $P$, we expect the agent can be aware of the complex step $t$, and initiate a human intervention. After providing precise guidance $g_{t}$ from a human or advanced model $\mathcal{F}_{s}$, the agent can arrive next step $t+1$, formed by $\mathcal{F}_{s}(a_{t}^{s}|s_{t},h_{t-1},\tau_{i},o_{t})$.
@@ 3.2 Challenge of Over-execution
To investigate the performance of existing GUI agents, we randomly select 350 instructions from three complex scenarios (Figure 9 and 10) to evaluate them. Then, we select the autonomous GUI agents: Qwen2-VL-7B and OS-Atlas-Pro-7B, and the interactive GUI agent assisted at each step by $\mathtt{GPT}$-$\mathtt{4o}$ in our pilot evaluation. Following the setting of Zhang and Zhang (2024) and Wu et al. (2024b), we report their performance in terms of action-Type success rate, the step-wise success rate (SR), and task success rate (TSR).
@@ 4.1 Collaborative Probing Framework
Confidence Annotation. Our confidence probing framework employs an agent-critic collaborative paradigm. To address the challenge of dynamic evaluation and expand coverage to commercial applications, the framework first utilizes Android Studio to connect real mobile devices and establish bidirectional communication with the probed GUI agent $\mathcal{F}_{p}$ deployed at the service station. Second, inspired by Ma et al. (2024b); Wang et al. (2024a), we assume that the layout-parse model enhanced $\mathtt{GPT}$-$\mathtt{4o}$ is the state-of-the-art critic model $\mathcal{F}_{c}$, capable of effectively supervising and guiding $\mathcal{F}_{p}$, thereby ensuring the dynamic probing of the entire trajectory. Additionally, $\mathcal{F}_{c}$ can monitor the entire probing process, including the planned schedule, current step, and instruction completion. The details of the prompt are provided in Appendix C.1.
Specifically, given a user instruction $\tau_{i}$, $\mathcal{F}_{p}$ predicts the next action $a_{t}^{p}$ using the action prompt $P_{p}$ at step $t$, followed by $\mathcal{F}_{p}(a_{t}^{p}|P_{p}(s_{t},h_{t-1},\tau_{i}))$. For example, $\mathcal{F}_{p}$ responds to the instruction $\tau_{i}$ with the first step “Open Amazon APP” as “CLICK <616, 371>”. Meanwhile, $\mathcal{F}_{c}$ first generates a plan schedule $L$ using the prompt $P_{l}$. Based on the current step $L_{t}$ at step $t$, it will also respond with a supervisory action $a_{t}^{c}$ using the action prompt $P_{c}$. Subsequently, $\mathcal{F}_{c}$ evaluates the effectiveness of the current step execution with the scoring prompt $P_{h}$:
$f_{\text{score}}:<\tau_{i},L_{t},s_{t},a_{t}^{p},a_{t}^{c},h_{t-1}>\stackrel{{\scriptstyle\mathcal{F}_{c}}}{{\rightarrow}}\text{score}_{t},$ (1)
@@ 4.2 Confidence-driven Interaction
Confidence Scoring Integration. Employing the trajectory from the collaborative probing framework, we introduce OS-Kairos, which integrates confidence scoring with the probed GUI agent $\mathcal{F}_{p}$. Specifically, we employ supervised training to fine-tune $\mathcal{F}_{p}$. Formally, the training objective $\mathcal{L}_{\text{{OS-Kairos}}}$ of the next-word prediction can be expressed as:
@@
OS-Kairos 100.0033.33↑ 100.00100.00↑ 100.0080.00↑ 99.852.05↑ 96.3333.87↑ 100.000.71↑ 92.8629.74↑ 99.883.98↑ 95.9034.54↑ 88.2073.91↑
Table 2: Comparison of OS-Kairos with baselines in complex scenarios (zero-shot setting). We report the overall accuracy for Type, SR, and TSR, along with fine-grained accuracy for each action. Subscripts indicate relative improvement over the OS-Atlas-Pro-7B, with the best result highlighted in bold.
@@
where human intervention is triggered if the current action’s confidence is below $\gamma$, otherwise, it is automatic. Notably, OS-Kairos switches to autonomous mode if the $\gamma$ is set to minimum value, or to fully interactive GUI if it is set to maximum value.
@@ 5.1 Experiment Setup
Datasets. Thanks to the confidence probing framework, we can evaluate the OS-Kairos in complex scenarios by splitting the generated trajectories. Moreover, we evaluate it on established benchmarks such as AITZ and Meta-GUI. Details are provided in Appendix C.2.
Models. In the confidence probing framework, we use the open-source GUI agent OS-Atlas-Pro-7B Wu et al. (2024b) as our probing model. Our objective is to probe the confidence score of the GUI model at each step, thereby introducing OS-Kairos to enhance its effectiveness. Additionally, we use $\mathtt{GPT}$-$\mathtt{4o}$ Achiam et al. (2023) as our critic model. The layout-parse model is resnet18 and convnextTiny for OCR-detection and recognition models, respectively Tang et al. (2019).
@@
$\bullet$ Multimodal API-based models. We consider MLLM-powered GUI agents, including $\mathtt{GPT}$-$\mathtt{4o}$ Achiam et al. (2023), $\mathtt{GLM}$-$\mathtt{4V}$-$\mathtt{Plus}$ GLM et al. (2024) and $\mathtt{Qwen}$-$\mathtt{VL}$-$\mathtt{MAX}$ Bai et al. (2023), which are strong baselines in zero-shot settings.
$\bullet$ Multimodal Open-source models. In the zero-shot setting, we also consider GUI-adapted MLLMs, including CogAgent Hong et al. (2024), Auto-UI Zhang and Zhang (2024), Qwen2-VL-7B Bai et al. (2023), OS-Atlas-Pro-7B Wu et al. (2024b). In the fine-tuning setting, we compare OS-Kairos with fine-tuned models on datasets.
Metrics. Following Wu et al. (2024b), we report the action type accuracy (Type), step-wise success rate (SR), and task success rate (TSR). Besides, we evaluate the human intervention success rate (HSR), intervention precision (IP), autonomous precision (AP), and relative efficiency (RE). More details of metrics and implementation can be found in Appendix C.3 and Appendix C.4.
@@
OS-Kairos ZS 99.3683.33↑ 100.00100.00↑ 94.7394.73↑ 99.815.28↑ 96.6659.37↑ 98.8338.60↑ 95.3280.12↑ 98.4932.40↑ 96.3672.77↑ 87.7187.29↑
Table 3: Comparison of OS-Kairos in the fine-tuning setting. ZS and FT denote zero-shot and fine-tuning evaluations, respectively. We report overall accuracy for Type, SR, and TSR, as well as fine-grained accuracy for each action. Subscripts indicate relative improvement over the ZS baseline, with the best result highlighted in bold.
@@ 5.2 Main Results
We present comparison results for complex scenarios and two benchmarks with zero-shot settings in Table  3, Appendix  C.6.1, and Appendix C.6.2. Table2 provides a comprehensive comparison of fine-tuning settings. Our key findings are as follows: In zero-shot setting: Superior Performance and Better Effectiveness. Without changing the model capabilities, OS-Kairos significantly outperforms the zero-shot baseline across three datasets, highlighting its effectiveness. Specifically, the adaptive interaction of OS-Kairos effectively identifies complex steps that trigger human intervention. This not only improves the prediction accuracy for each action but also enhances overall performance. For example, it achieves 95.90% in SR and 88.20% in TSR for complex scenarios. Although API-based and proprietary models realizes domain enhancement for GUI tasks, they cannot identify complex steps, resulting in over-execution and task failure. Moreover, OS-Kairos yields promising results on the other two datasets when applying confidence scoring integration to the original dataset, highlighting its generality (see Appendix C.6.3).
In fine-tuning setting: Competitive Performance and Precise Improvement. Although fine-tuning can alleviate the over-execution of GUI agents, OS-Kairos still outperforms them, achieving high SR and notable improvements in TSR. For example, it shows relative improvements ranging from 26.09% to 85.72% in complex scenarios. Furthermore, OS-Kairos achieves precise improvements by identifying complex steps (e.g., $\mathtt{SCROLL}$), while fine-tuning may introduce side effects in specific actions and encounter optimization bottlenecks.
@@
Confidence Scoring: Effective interaction. In Figure 4, OS-Kairos shows accurate confidence evaluation (HSR). Thus, It does not interfere with the autonomous steps (AP), as seen in complex scenarios (96.44%) and MetaGUI (93.18%). Notably, OS-Kairos achieves over 70% precision in all human intervention steps, highlighting its effective interaction. With high-quality sampling, we consider OS-Kairos’s precision can be improved further (e.g., AITZ). Additionally, the ablation study of the critic model shows that $\mathtt{GPT}$-$\mathtt{4o}$ is the optimal choice (see Appendix C.7).

### v2

# Zhuosheng Zhang1, Gongshen Liu1 1

@@ Abstract
Pengzhou Cheng1thanks:   Equal contribution. Corresponding authors. This work is partially supported by the Joint Funds of the National Natural Science Foundation of China (U21B2020), National Natural Science Foundation of China (62406188), and Natural Science Foundation of Shanghai (24ZR1440300)., Zheng Wu1, Zongru Wu1, Tianjie Ju1, Aston Zhang2, Zhuosheng Zhang1, Gongshen Liu1 1School of Computer Science, Shanghai Jiao Tong University 2GenAI, Meta {cpztsm520,wzh815918208,wuzongru,jometeorie,zhangzs,lgshen}@sjtu.edu.cn aston@meta.com
@@ 1 Introduction
To annotate the confidence scores for the probed GUI agents in real-world scenarios, we first design a collaborative confidence probing framework. Inspired by Chen et al. (2024a), this framework integrates a layout parsing model Tang et al. (2019) and the most capable proprietary model, GPT-4o Achiam et al. (2023) to function as a critic model. The critic model is used to supervise plan scheduling and confidence score based on our curated instructions to address complex scenarios. This framework is the first toolkit designed to identify when human intervention is necessary, generate confidence scores, and facilitate the automated construction of GUI trajectories. To further integrate confidence scoring into the probed GUI agent, we validate and refine these GUI trajectories and then fine-tune the model. This approach ensures action prediction accuracy while improving adaptability of human intervention.
@@ 2.1 MLLM-powered GUI Agents
Early efforts to build GUI agents rely on the availability of commercial MLLMs. These agents can be built through prompt learning based on GPT-4o or Gemini-Pro Vision, e.g., AppAgent Zhang et al. (2023) and Mobile-Agent Wang et al. (2024a). However, practitioners are concerned about the costs associated with API requests and the delays in inference on mobile devices. Recent studies have focused on fine-tuning to optimize foundation models. On the one hand, they work on performing fine-grained visual understanding Bai et al. (2023), model scaling laws Chen et al. (2024b), multimodal information integration Hong et al. (2024), and GUI grounding enhancements Wu et al. (2024b); Qin et al. (2025) in the pre-training phase. On the other hand, researchers fine-tune the foundation model on GUI-specific datasets to enhance action orientation Wu et al. (2024a), planning decision Zhang et al. (2024c), perception enhancement Ma et al. (2024b), and reasoning Zhang and Zhang (2024); Zhang et al. (2024b). Moreover, a framework based on reinforcement learning (RL) designed specifically for the GUI agents can further enhance robustness Zhou et al. (2024); Liu et al. (2024); Wang et al. (2024b).
@@ 2.2 Capability Probing for GUI Agent
GUI agent-oriented capability probing is critical for real-world applications Deka et al. (2017). Generally, the capability of GUI agents can be probed by releasing benchmark datasets. Examples like UIBert Bai et al. (2021), SeeClick Cheng et al. (2024), and OS-Copilot Wu et al. (2024b), which investigate the problem of grounding understanding to UI elements on a screen. Besides, large-scale, diverse, and high-quality trajectory datasets can identify challenges of action prediction in terms of effectiveness (e.g., PixelHelp Li et al. (2020), Meta-GUI Sun et al. (2022), and AndroidWorld Rawles et al. (2024)), task complexity (e.g., Mobile-Bench Deng et al. (2024) and GUI Odyssey Lu et al. (2024)), and data-scaling (e.g., AITW Rawles et al. (2024) and AndroidControl Li et al. (2024)). After identifying the capacity bottleneck of GUI agents, the introduction of specific strategies (e.g., planning lists Zhang et al. (2024c), action chains Zhang and Zhang (2024); Zhang et al. (2024b), and supplementary data) further enhance the environment perception. However, most benchmark datasets rely on crowdsourcing and human annotation.
@@ 3 Pilot Experiments
In this section, we first define GUI agent paradigms and then investigate the over-execution issue of the existing GUI agent in three complex scenarios. As shown in Figure 2, GUI agents confront substantial risks in real-world scenarios, such as those involving ambiguous user instructions (e.g., information absence and account logout), unexpected interruptions (e.g., hallucinations and shortcuts), and environmental hijacks (e.g., Pop-up hijacking and permission unauthorized).
@@ 3.1 GUI Agent Paradigm
Interactive GUI Agent. Given an interactive GUI agent $\mathcal{F}_{i}$ and system prompt $P$, we expect the agent can be aware of the complex step $t$, and initiate a human intervention. After providing precise guidance $g_{t}$ from a human or advanced model $\mathcal{F}_{s}$, the agent can arrive at the next step $t+1$, formed by $\mathcal{F}_{s}(a_{t}^{s}|s_{t},h_{t-1},\tau_{i},o_{t})$.
@@ 3.2 Challenge of Over-execution
To investigate the performance of existing GUI agents, we randomly select 350 instructions from three complex scenarios (Figures 9 and 10) to evaluate them. Then, we select the autonomous GUI agents: Qwen2-VL-7B and OS-Atlas-Pro-7B, and the interactive GUI agent assisted at each step by GPT-4o in our pilot evaluation. Following the setting of Zhang and Zhang (2024) and Wu et al. (2024b), we report their performance in terms of action-Type success rate, the step-wise success rate (SR), and task success rate (TSR).
@@ 4.1 Collaborative Probing Framework
Confidence Annotation. Our confidence probing framework employs an agent-critic collaborative paradigm. To address the challenge of dynamic evaluation and expand coverage to commercial applications, the framework first utilizes Android Studio to connect real mobile devices and establish bidirectional communication with the probed GUI agent $\mathcal{F}_{p}$ deployed at the service station. Second, inspired by Ma et al. (2024b); Wang et al. (2024a), we assume that the layout-parse model enhanced GPT-4o is the state-of-the-art critic model $\mathcal{F}_{c}$, capable of effectively supervising and guiding $\mathcal{F}_{p}$, thereby ensuring the dynamic probing of the entire trajectory. Additionally, $\mathcal{F}_{c}$ can monitor the entire probing process, including the planned schedule, current step, and instruction completion. The details of the prompt are provided in Appendix C.1.
Specifically, given a user instruction $\tau_{i}$, $\mathcal{F}_{p}$ predicts the next action $a_{t}^{p}$ using the action prompt $P_{p}$ at step $t$, followed by $\mathcal{F}_{p}(a_{t}^{p}|P_{p}(s_{t},h_{t-1},\tau_{i},o_{t}))$. For example, $\mathcal{F}_{p}$ responds to the instruction $\tau_{i}$ with the first step “Open Amazon APP” as “CLICK <616, 371>”. Meanwhile, $\mathcal{F}_{c}$ first generates a plan schedule $L$ using the prompt $P_{l}$. Based on the current step $L_{t}$ at step $t$, it will also respond with a supervisory action $a_{t}^{c}$ using the action prompt $P_{c}$. Subsequently, $\mathcal{F}_{c}$ evaluates the effectiveness of the current step execution with the scoring prompt $P_{h}$:
$f_{\text{score}}:<\tau_{i},$ $L_{t},s_{t},a_{t}^{p},a_{t}^{c},$ (1)
$h_{t-1},o_{t}>\stackrel{{\scriptstyle\mathcal{F}_{c}}}{{\rightarrow}}\text{score}_{t},$
@@ 4.2 Confidence-driven Interaction
Confidence Scoring Integration. Employing the trajectory from the collaborative probing framework, we introduce OS-Kairos, which integrates confidence scoring with the probed GUI agent $\mathcal{F}_{p}$. Specifically, we employ supervised training to fine-tune $\mathcal{F}_{p}$. Formally, the training objective $\mathcal{L}_{\text{\text{OS-Kairos}}}$ of the next-word prediction can be expressed as:
@@
OS-Kairos 100.033.33↑ 100.0100.0↑ 100.080.00↑ 99.852.05↑ 96.3333.87↑ 100.00.71↑ 92.8629.74↑ 99.883.98↑ 95.9034.54↑ 88.2073.91↑
Table 2: Comparison of OS-Kairos with baselines in complex scenarios (zero-shot setting). We report the overall accuracy for Type, SR, and TSR, along with fine-grained accuracy for each action. Subscripts indicate absolute improvement over the OS-Atlas-Pro-7B, with the best result highlighted in bold.
@@
where human intervention is triggered if the confidence score of current action falls below a threshold $\gamma$, otherwise, the OS-Kairos operates autonomously. It is noted that OS-Kairos switches to autonomous mode if $\gamma$ is set to the minimum value, or to fully interactive GUI if $\gamma$ is set to the maximum value.
@@ 5.1 Experiment Setup
Datasets. Thanks to the confidence probing framework, we can evaluate the OS-Kairos against complex scenarios by splitting the generated trajectories. Moreover, we evaluate it on established benchmarks such as AITZ and Meta-GUI. Details are provided in Appendix C.2.
Models. In the confidence probing framework, We use the open-source model OS-Atlas-Pro-7B Wu et al. (2024b) as our probing model. Our objective is to probe the confidence score of the GUI agent at each step, thereby introducing OS-Kairos to enhance its effectiveness. Additionally, we use GPT-4o Achiam et al. (2023) as our critic model. The layout-parse model is resnet18 and convnextTiny for OCR detection and recognition, respectively Tang et al. (2019).
@@
$\bullet$ Multimodal API-based models. We consider MLLM-powered GUI agents, including GPT-4o Achiam et al. (2023), GLM-4V-Plus GLM et al. (2024) and Qwen-VL-MAX Bai et al. (2023), which are strong baselines in zero-shot settings.
$\bullet$ Multimodal Open-source models. In the zero-shot setting, we include GUI-adapted MLLMs as our baseline methods, such as CogAgent Hong et al. (2024), Auto-UI Zhang and Zhang (2024), Qwen2-VL-7B Bai et al. (2023), and OS-Atlas-Pro-7B Wu et al. (2024b). For the fine-tuning setting, we compare OS-Kairos against models fine-tuned on the Kairos dataset as well as on two established GUI benchmarks.
Metrics. Following Wu et al. (2024b), we report the action type accuracy (Type), step-wise success rate (SR), and task success rate (TSR). Besides, we evaluate the human intervention success rate (HSR), intervention precision (IP), autonomous precision (AP), and relative efficiency (RE). More details of metrics can be found in Appendix C.3.
Implementation Details. For each dataset, we randomly split 80% trajectories as training data, and 20% trajectories as testing data. Dataset statistics are presented in Appendix C.2. To ensure a fair comparison with the baseline, we use GPT-4o to score between the probing model and ground truth actions in the two benchmarks, without relying on high-quality sampling. In the zero-shot scenario, we evaluate the GUI agent directly using prompt learning. In the fine-tuning scenario, we fine-tune the model for 8 epochs on the corresponding dataset with a learning rate of 1e-5. In the interactive mode, if not specifically mentioned, the threshold $\gamma$ is set to 4. When human intervention is required at the current step, OS-Kairos uses ground truth for the evaluation of the data set or human guidance for the dynamic evaluation.
@@
OS-Kairos ZS 99.3683.33↑ 100.0100.0↑ 94.7394.73↑ 99.815.28↑ 96.6659.37↑ 98.8338.60↑ 95.3280.12↑ 98.4932.40↑ 96.3672.77↑ 87.7187.29↑
Table 3: Comparison of OS-Kairos in the fine-tuning setting. ZS and FT denote zero-shot and fine-tuning evaluations, respectively. We report overall accuracy for Type, SR, and TSR, as well as fine-grained accuracy for each action. Subscripts indicate absolute improvement over the ZS baseline, with the best result highlighted in bold.
@@ 5.2 Main Results
We present comparison results for complex scenarios and two benchmarks with zero-shot settings in Table 2, Appendix C.5.1, and Appendix C.5.2. Table 3 provides a comprehensive comparison of fine-tuning settings. Our key findings are as follows: In zero-shot setting: Superior Performance and Better Effectiveness. Without changing the model capabilities, OS-Kairos significantly outperforms the zero-shot baseline in three datasets, highlighting its effectiveness. The adaptive interaction of OS-Kairos effectively identifies complex steps that trigger human intervention. This not only improves the prediction accuracy for each action, but also enhances overall performance. For example, it achieves 95.90% in SR and 88.20% in TSR for complex scenarios. Although API-based and proprietary models realize domain enhancement for GUI tasks, they cannot identify complex steps, resulting in over-execution and task failure. Moreover, OS-Kairos yields promising results on the other two datasets when applying confidence scoring integration to the original dataset, highlighting its generality (see Appendix C.5.3).
In fine-tuning setting: Competitive Performance and Precise Improvement. Although fine-tuning can alleviate over-execution of GUI agents, OS-Kairos still outperforms them, achieving high SR and notable improvements in TSR. For example, it shows absolute improvements ranging from 26.09% to 85.72% in complex scenarios. Furthermore, OS-Kairos achieves precise improvements by identifying complex steps (e.g. $\mathtt{SCROLL}$) while fine-tuning can introduce side effects in specific actions and encounter optimization bottlenecks.
@@
Confidence Scoring: Effective interaction. In Figure 4, OS-Kairos shows accurate confidence evaluation (HSR). Thus, it does not interfere with the autonomous steps (AP), as seen in complex scenarios (96.44%) and Meta-GUI (93.18%). Notably, OS-Kairos achieves over 70% precision in all human intervention steps. With high-quality sampling, we consider that OS-Kairos’s precision can be improved further (e.g., AITZ).